OpenStudy (jessiegonzales):

MEDAL!!!!! Find the area of the regular polygon. Give the answer to the nearest tenth. Hexagon with a radius of 5 in. A. 65.0 in^2 <------ B. 129.9 in^2 C. 259.8 in^2 D. 53.0 in^2

OpenStudy (stefrheart):

First draw the hexagon

OpenStudy (jessiegonzales):

Ok! Is my answer correct?

OpenStudy (stefrheart):

what is your answer?

OpenStudy (jessiegonzales):

A

OpenStudy (anonymous):

Area of Polygon = 1/2 x n x Radius^2 x sin(2 x pi/n) Area of Polygon = 1/2 × (6) x (5 in)^2 x sin(2 x pi/6) Area of Polygon = 3 x 25 x 0.86602540378 in^2 Area of Polygon = 64.9519052835 in^2 Area of Polygon = 65.0 in^2(in the nearest tenth.)

OpenStudy (jessiegonzales):

Anyone know?

OpenStudy (anonymous):

the answer is A

OpenStudy (jessiegonzales):

Thanks

OpenStudy (anonymous):

Area of Polygon = 1/2 x n x Radius^2 x sin(2 x pi/n) Area of Polygon = 1/2 x (6) x (5 in)^2 x sin(2 x pi/6) Area of Polygon = 3 x 25 x 0.86602540378 in^2 Area of Polygon = 64.9519052835 in^2 Area of Polygon = 65.0 in^2(in the nearest tenth.)

OpenStudy (anonymous):

you are welcome :)